# Exploring the Abilities of Large Language Models to Solve Proportional Analogies via Knowledge-Enhanced Prompting

### v1

@@ Abstract
Making analogies is fundamental to cognition. Proportional analogies, which consist of four terms, are often used to assess linguistic and cognitive abilities. For instance, completing analogies like “Oxygen is to Gas as <blank> is to <blank>” requires identifying the semantic relationship (e.g., “type of”) between the first pair of terms (“Oxygen” and “Gas”) and finding a second pair that shares the same relationship (e.g., “Aluminum” and “Metal”). In this work, we introduce a 15K Multiple-Choice Question Answering (MCQA) dataset for proportional analogy completion and evaluate the performance of contemporary Large Language Models (LLMs) in various knowledge-enhanced prompt settings. Specifically, we augment prompts with three types of knowledge: exemplar, structured, and targeted. Our results show that despite extensive training data, solving proportional analogies remains challenging for current LLMs, with the best model achieving an accuracy of 55%. Notably, we find that providing targeted knowledge can better assist models in completing proportional analogies compared to providing exemplars or collections of structured knowledge.
@@
Exploring the Abilities of Large Language Models to Solve
Proportional Analogies via Knowledge-Enhanced Prompting
Thilini Wijesiriwardene1, Ruwan Wickramarachchi1, Sreeram Vennam2, Vinija Jain4,5thanks:  Work does not relate to position at Meta., Aman Chadha3,5thanks:   Work does not relate to position at Amazon., Amitava Das1, Ponnurangam Kumaraguru2, Amit Sheth 1, 1AI Institute, University of South Carolina, USA, 2IIIT Hyderabad, India 3Amazon GenAI, USA, 4Meta, USA 5Stanford University, USA Correspondence: thilini@sc.edu
@@ 1 Introduction
Our primary contribution lies in conducting a comprehensive evaluation of nine GenAI models, specifically assessing their performance in solving proportional analogies presented in a multiple-choice format. Considering the limitations of existing proportional analogy datasets, which typically comprise fewer than a thousand data points and a restricted range of relation types, we present a substantially larger dataset. Our dataset contains 15K proportional analogies with 238 distinct relation types. We evaluate the nine GenAI models on the 15K dataset using four distinct prompting techniques: (i) Zero-shot Prompting, where no additional knowledge is incorporated into the prompt, (ii) Few-shot Prompting, where exemplar knowledge in the form of examples from the dataset is included in the prompt, (iii) Structured Knowledge Prompting (SKP), where the prompt is augmented with structured knowledge in the forms of lexical, commonsense, and world knowledge drawn from WordNet McCrae et al. (2019), ConceptNet Speer et al. (2017), and Wikidata Vrandečić and Krötzsch (2014) respectively and (iv) Targeted Knowledge Prompting (TKP), which integrates targeted knowledge in the form of specific semantic relationships necessary for solving proportional analogies, along with the cognitive process behind such reasoning. To the best of our knowledge, this study is the first to explore knowledge-enhanced prompting strategies for solving proportional analogies.
@@ 3.1 Dataset Creation
We introduce a 15K dataset of proportional analogies containing 5-way, 4-way and 3-way MCQs. Table 1 presents the dataset statistics along with examples from the dataset. We generate 14K questions out of the 15K based on the work by Yuan et al. (2023). Yuan et al. (2023) introduced a an automatically generated million-scale analogy knowledge base with diverse relational structures among the analogous term pairs. We adopt this resource to develop n-way (n=[3, 4, 5]) MCQs as follows. A single n-way MCQ consist of a pair of terms representing the question and five term pairs representing the answer choices, among which only one term pair is the correct answer. The semantic relationship between the term pair in the question is the same as the semantic relationship shared between the term pair which is the correct answer. The rest of the incorrect answer choices consist of term pairs with different semantic relationships among them.
Thousand data points out of the 15K are borrowed from work by Ushio et al. (2021); Turney and Littman (2003); Boteanu and Chernova (2015) [^3] Unlike the 14K MCQs created based on AnalogyKB, these 1K data points do not provide the semantic relationship shared between the question term pair explicitly, therefore we employ two NLP researchers to discuss and manually identify the shared semantic relationship. and contain 5-way, 4-way and 3-way MCQs. We highlight that, compared to previous proportional analogy MCQ datasets used for research Webb et al. (2023); Ushio et al. (2021), the current dataset provides a significant increase in question quantity ($\sim$15-times) and diversity (with regard to diversity of semantic relations among terms). Our dataset also includes the semantic relationship shared by the question term pair compared to other datasets that do not include this information Ushio et al. (2021); Turney and Littman (2003); Boteanu and Chernova (2015).
@@
Amount 14386 610 4    Total # relation types 238
Table 1: Dataset statistics. The dataset consist of 15K MCQs that share 238 semantic relation types among them.
@@ 3.3 Prompting Techniques
Currently, the most popular approach to Multiple Choice Question Answering (MCQA) is via cloze-style prompting Brown (2020); Robinson et al. (2023) where each answer choice is concatenated to the question separately and scored independently by the language model (LM). This style of prompting is problematic since it prevents the LM from comparing and contrasting all available options simultaneously. Additionally, it is computationally expensive, as it requires multiple forward passes through the LM to identify the correct answer Robinson et al. (2023). To address these limitations, we adopt the prompt phrasing introduced by Robinson et al. (2023) with task-specific modifications. Specifically, the question and its symbol-enumerated candidate answers are provided to the model as a single prompt. Robinson et al. (2023) do not include specific instructions in the prompt for the model to output only the choice symbol. But we observe that adding such specific instructions reduce the model hallucinations. Therefore we use specific, non-ambiguous language to instruct the model to only output the relevant choice symbol. The prompting techniques are detailed below (See example prompts in appendix C).
@@ 4 Experimental Setting
We have conducted a comprehensive set of experiments across nine GenAI models over six prompt variants on a 15K dataset, totalling to 54 ($69X$ 6) experiments.
4.1 Implementation Details
We use API requests for GPT-3.5-Turbo and checkpoints from Hugging face [^9] https://huggingface.co/models for open-source models. The models are evaluated with following hyper parameter settings, temperature = 0.1, top_p=0.1 and repetition_penalty=1.2 to elicit more concrete answers for the MCQs. We use Sentence Transformers [^10] https://sbert.net/ to identify semantically similar exemplars and to perform semantic knowledge filtering. We utilize Wikidata knowledge from Wang et al. (2021b), ConceptNet knowledge from conceptnet5111111 https://github.com/commonsense/conceptnet5/wiki/Downloads and WordNet knowledge from Open English WordNet (2023) [^12] https://github.com/globalwordnet/english-wordnet?tab=readme-ov-file.
@@ 5.1 Model Performance and Prompting Techniques
The highest overall performance was attained by GPT-3.5-Turbo, achieving an EMA of 55.25%. This result underscores the challenge that proportional analogies pose for current state-of-the-art GenAI models. This accuracy was obtained through Targeted Knowledge Prompting where the prompt was enhanced with targeted knowledge. Interestingly, the same model, when enhanced with structured knowledge, underperformed with an accuracy of  38% (EMA for SKP[random] is 38.29% and SKP[semantic] is 38.79%), compared to Zero-shot prompting (EMA 45.7%). This suggests that simply adding knowledge, even from diverse sources, may not be beneficial for cognitively demanding tasks such as proportional analogy completion. Out of the nine models four (Falcon, Flan, Mistral and GPT-3.5-Turbo) performs the best when prompted with Targeted Knowledge Prompts and two (GPT2 and Orca) performs the best with Zero-shot prompts with no knowledge enhancement. CodeT5 performs the best with one-shot prompts and Zephyr and CodeParrot performs the best with five-shot prompts. We also observe that models trained specifically on code generation such as CodeT5 and CodeParrot (specially CodeParrot) perform at the lower end of the spectrum despite the demonstrated abilities of them to perform well on other MCQ datasets Robinson et al. (2023). We believe this is due to the challenging nature of the proportional analogy completion task.
@@ 5.2 Role of Structured Knowledge in Model Performance
Although enhancing prompts with structured knowledge does not consistently improve model performance compared to other prompting techniques, SKP[semantic] leads to slight increases in EMA values (ranging from 0.01% to 1.32%) compared to SKP[random], across all models except GPT-2 and Mistral (see Table 4). We identified a subset of MCQs (19.96%) where all three types of knowledge were available and conducted additional experiments to evaluate the individual contribution of each knowledge type to EMA (we employed SKP[semantic ] prompting). Our results show (See table 3 in Appendix B) that incorporating each of the three knowledge types separately into prompts leads to very similar EMA values (when averaged across all nine models). Specifically, prompts enhanced only with Wikidata knowledge resulted in an average EMA of 14.57%, while using only WordNet or only ConceptNet yielded average EMAs of 14.41% and 14.34%, respectively.
@@ 6 Conclusion and Future Work
We explore the ability of nine LLMs to solve proportional analogies via different knowledge-enhanced prompting techniques on a 15K MCQ dataset that we introduce. Our extensive experiments show that most LLMs perform their best when targeted knowledge is incorporated into prompts as opposed to exemplar knowledge and structured knowledge.
Even though the several models used in the current study are instruction finetuned versions of their base models, they are not necessarily finetuned for the proportional analogy completion task. Therefore, there is room to explore the performance of models specifically finetuned for proportional analogy completion. Our study currently evaluate manual prompting techniques which are known to be brittle. We believe that evaluating automatic prompting approaches for the same task would be compelling.
@@ 7 Limitations
When incorporating knowledge paths in SKP, the exact semantic relationship between the question term pair may sometimes be retrieved from the knowledge sources. We define this specific information as targeted knowledge. As a result, these instances can be categorized as SKP with targeted knowledge. However, we do not currently verify or adjust for them.
In this work, we employed manual prompt creation, where slight variations can lead to significant differences in model outputs Zhao et al. (2021). We did not account for this variability in the current study by experimenting with several prompt templates for each prompting technique.
@@ Appendix A Model Details
Model Name WD knowledge only CN knowledge only WN knowledge only All knowledge available
Falcon 25.14 25.38 25.04 24.14
FlanT5 19.63 28.15 30.7 12.19
GPT2 4.11 2.4 1 5.84
Mistral 22.84 22.07 23.34 24.11
Orca 14.62 9.52 9.28 22.24
Zephyr 16.16 12.49 9.75 14.76
CodeT5 21.47 21.64 22.7 20.7
CodeParrot 0 0 0 0
GPT-3.5-Turbo 7.13 7.43 7.9 7.2
@@
Figure 5: Prompt Lengths vs. Peak Performance
Appendix B Performance and Additional Results
B.1 Model Performance vs. Prompt Length (PL)
@@ Appendix B Performance and Additional Results
Appendix C Prompts
Figues 6, 7, 8, 9 and 10 illustrates example prompts provided to models.

### v2

# KnowledgePrompts: Exploring the Abilities of Large Language Models to Solve Proportional Analogies via Knowledge-Enhanced Prompting

@@ Abstract
Making analogies is fundamental to cognition. Proportional analogies, which consist of four terms, are often used to assess linguistic and cognitive abilities. For instance, completing analogies like “Oxygen is to Gas as <blank> is to <blank>” requires identifying the semantic relationship (e.g., “type of”) between the first pair of terms (“Oxygen” and “Gas”) and finding a second pair that shares the same relationship (e.g., “Aluminum” and “Metal”). In this work, we introduce a 15K Multiple-Choice Question Answering (MCQA) dataset for proportional analogy completion and evaluate the performance of contemporary Large Language Models (LLMs) in various knowledge-enhanced prompt settings. Specifically, we augment prompts with three types of knowledge: exemplar, structured, and targeted. Our results show that despite extensive training data, solving proportional analogies remains challenging for current LLMs, with the best model achieving an accuracy of 55%. Notably, we find that providing targeted knowledge can better assist models in completing proportional analogies compared to providing exemplars or collections of structured knowledge. Our code and data are available at: https://github.com/Thiliniiw/KnowledgePrompts/
@@
KnowledgePrompts: Exploring the Abilities of Large Language Models
to Solve Proportional Analogies via Knowledge-Enhanced Prompting
Thilini Wijesiriwardene1, Ruwan Wickramarachchi1, Sreeram Vennam2, Vinija Jain4thanks: Work does not relate to position at Meta., Aman Chadha3thanks: Work does not relate to position at Amazon., Amitava Das1, Ponnurangam Kumaraguru2, Amit Sheth1 1AI Institute, University of South Carolina, USA, 2IIIT Hyderabad, India 3Amazon GenAI, USA, 4Meta AI, USA Correspondence: thilini@sc.edu
@@ 1 Introduction
Our primary contribution lies in conducting a comprehensive evaluation of nine GenAI models, specifically assessing their performance in solving proportional analogies presented in a multiple-choice format. Considering the limitations of existing proportional analogy datasets, which typically comprise fewer than a thousand data points and a restricted range of relation types, we present a substantially larger dataset. Our dataset contains 15K proportional analogies with 236 distinct relation types. We evaluate the nine GenAI models on the 15K dataset using four distinct prompting techniques: (i) Zero-shot Prompting, where no additional knowledge is incorporated into the prompt, (ii) Few-shot Prompting, where exemplar knowledge in the form of examples from the dataset is included in the prompt, (iii) Structured Knowledge Prompting (SKP), where the prompt is augmented with structured knowledge in the forms of lexical, commonsense, and world knowledge drawn from WordNet McCrae et al. (2019), ConceptNet Speer et al. (2017), and Wikidata Vrandečić and Krötzsch (2014) respectively and (iv) Targeted Knowledge Prompting (TKP), which integrates targeted knowledge in the form of specific semantic relationships necessary for solving proportional analogies, along with the cognitive process behind such reasoning. To the best of our knowledge, this study is the first to explore knowledge-enhanced prompting strategies for solving proportional analogies.
@@ 3.1 Dataset Creation
We introduce a 15K dataset of proportional analogies containing 5-way, 4-way and 3-way MCQs. Table 1 presents the dataset statistics along with examples from the dataset. We generate 14K questions out of the 15K based on the work by Yuan et al. (2023). Yuan et al. (2023) introduced an automatically generated million-scale analogy knowledge base based on ConcepNet and Wikidata knowledge graphs. Yuan et al. (2023) acquire analogies of the same relations directly utilizing the concept pairs in the above-mentioned knowledge graphs. To acquire analogies of analogous relations (analogies consisting of two concept pairs with two relations that are analogous to each other) Yuan et al. (2023) utilize the in-context learning abilities of LLMs. We adopt this resource (specifically the analogies of same relations) to develop n-way (n=[3, 4, 5]) MCQs as follows. A single n-way MCQ consist of a pair of terms representing the question and n term pairs representing the answer choices, among which only one term pair is the correct answer. The semantic relationship between the term pair in the question is the same as the semantic relationship shared between the term pair which is the correct answer. The rest of the incorrect answer choices consist of term pairs with different semantic relationships among them.
Thousand data points out of the 15K are borrowed from work by Ushio et al. (2021); Turney and Littman (2003); Boteanu and Chernova (2015) [^3] Unlike the 14K MCQs created based on AnalogyKB, these 1K data points do not provide the semantic relationship shared between the question term pair explicitly, therefore we employ two NLP researchers to discuss and manually identify the shared semantic relationship. and contain 5-way, 4-way and 3-way MCQs. We highlight that, compared to previous proportional analogy MCQ datasets used for research Webb et al. (2023); Ushio et al. (2021), the current dataset provides a significant increase in question quantity ($\sim$15-times) and diversity (with respect to the diversity of semantic relations between terms). Our dataset also includes the semantic relationship shared by the question term pair compared to other datasets that do not include this information Ushio et al. (2021); Turney and Littman (2003); Boteanu and Chernova (2015). Our dataset contains 59 semantic relationship types with more than 10 instances each. The distribution of these relationships (focusing on the top 59 types) is depicted in Figure 2.
Figure 2: Distribution of Semantic relations. The distribution of the top 59 semantic relations (these are the frequencies of semantic relations between the question word pair )
@@
Amount 14386 610 4    Total # relation types 236
Table 1: Dataset statistics. The dataset consist of 15K MCQs that share 236 semantic relation types among them.
@@ 3.3 Prompting Techniques
Currently, the most popular approach to Multiple Choice Question Answering (MCQA) is via cloze-style prompting Brown (2020); Robinson et al. (2023) where each answer choice is concatenated to the question separately and scored independently by the language model (LM). This style of prompting is problematic since it prevents the LM from comparing and contrasting all available options simultaneously. Additionally, it is computationally expensive, as it requires multiple forward passes through the LM to identify the correct answer Robinson et al. (2023). To address these limitations, we adopt the prompt phrasing introduced by Robinson et al. (2023) with task-specific modifications. Specifically, the question and its symbol-enumerated candidate answers are provided to the model as a single prompt. Robinson et al. (2023) do not include specific instructions in the prompt for the model to output only the choice symbol. But we observe that adding such specific instructions reduce the model hallucinations. Therefore we use specific, non-ambiguous language to instruct the model to only output the relevant choice symbol. The prompting techniques are detailed below (See example prompts in appendix D).
@@ 4 Experimental Setting
We have conducted a comprehensive set of experiments across nine GenAI models over six prompt variants on a 15K dataset, totalling to 54 ($69X$ 6) experiments. The implementation details are included in Appendix B
@@ 5.1 Model Performance and Prompting Techniques
The highest overall performance was attained by GPT-3.5-Turbo, achieving an EMA of 55.25%. This result underscores the challenge that proportional analogies pose for current state-of-the-art GenAI models. This accuracy was obtained through Targeted Knowledge Prompting where the prompt was enhanced with targeted knowledge (See Figure 5). Interestingly, the same model, when enhanced with structured knowledge, underperformed with an accuracy of  38% (EMA for SKP[random] is 38.29% and SKP[semantic] is 38.79%), compared to Zero-shot prompting (EMA 45.7%). This suggests that simply adding knowledge, even from diverse sources, may not be beneficial for cognitively demanding tasks such as proportional analogy completion. Out of the nine models four (Falcon, Flan, Mistral and GPT-3.5-Turbo) performs the best when prompted with Targeted Knowledge Prompts and two (GPT2 and Orca) performs the best with Zero-shot prompts with no knowledge enhancement. CodeT5 performs the best with one-shot prompts and Zephyr and CodeParrot performs the best with five-shot prompts. We also observe that models trained specifically on code generation such as CodeT5 and CodeParrot (specially CodeParrot) perform at the lower end of the spectrum despite the demonstrated abilities of them to perform well on other MCQ datasets Robinson et al. (2023). We believe this is due to the challenging nature of the proportional analogy completion task.
@@ 5.2 Role of Structured Knowledge in Model Performance
Although enhancing prompts with structured knowledge does not consistently improve model performance compared to other prompting techniques, SKP[semantic] leads to slight increases in EMA values (ranging from 0.01% to 1.32%) compared to SKP[random], across all models except GPT-2 and Mistral (see Table 2). We identified a subset of MCQs (19.96%) where all three types of knowledge were available and conducted additional experiments to evaluate the individual contribution of each knowledge type to EMA (we employed SKP[semantic ] prompting). Our results show (see Figure 4, for complete results, see table 4 in Appendix C) that incorporating each of the three knowledge types separately into prompts leads to very similar EMA values (when averaged across all nine models). Specifically, prompts enhanced only with Wikidata knowledge resulted in an average EMA of 14.57%, while using only WordNet or only ConceptNet yielded average EMAs of 14.41% and 14.34%, respectively.
@@
EMA % of Large Language Model
Semantic Relationship Type Falcon FlanT5 GPT2 Mistral Orca Zephyr CodeT5 CodeParrot GPT-3.5-Turbo
part of 23.50 25.05 20.19 24.95 23.98 9.03 13.88 0.00 27.18
is a 25.78 29.20 20.09 25.50 22.08 6.84 19.37 0.00 31.34
at location 29.34 31.85 21.43 27.80 19.31 11.20 18.73 0.39 32.24
follows 22.07 28.46 15.43 22.34 23.94 3.19 14.36 0.27 35.11
producer 25.94 37.97 22.19 24.87 23.80 7.22 13.90 0.00 53.21
[…]
Table 3: Performance of each LLM across semantic relations. We report the performance of each LLM on the top five semantic relations in the dataset. For each LLM, the relation with the highest performance is highlighted in green, while the relation with the lowest performance is highlighted in orange. Additionally, the average performance across all five relation types is calculated and highlighted in grey.
@@
5.5 Diversity of Semantic Relationships vs. Model Performance
As elaborated in Section 3.1, our dataset encompasses 236 unique semantic relation types, with the frequencies of the top five relations detailed in Table 1. To further elucidate the performance of each LLM, we assess their results on these top five semantic relations using targeted knowledge prompts (refer to Table 3). Consistent with prior findings, GPT-3.5-Turbo achieves the highest average performance across the top five relations, followed by FlanT5. For both models, the MCQs involving the "part of" relation pose the greatest challenge, whereas the "producer" relation is the easiest to solve. Similarly, across all nine LLMs, the "part of" and "follows" relations emerge as the most difficult, while the "at location" relation proves to be the easiest.
@@ 6 Conclusion and Future Work
We evaluate nine LLMs on a 15K MCQ dataset to assess their ability to solve proportional analogies using various knowledge-enhanced prompting techniques. Our experiments reveal that LLMs perform best when targeted knowledge is integrated into prompts, outperforming exemplar and structured knowledge.
While several of the models used are instruction-finetuned versions of their base models, they are not specifically finetuned for proportional analogy completion, leaving room for improvement. Additionally, our study focuses on manual prompting techniques, which are brittle; exploring automatic prompting approaches could yield more robust results.
@@ 7 Limitations
In SKP, knowledge paths may occasionally provide the exact semantic relationship between question term pairs, defined as targeted knowledge. These instances can be classified as SKP with targeted knowledge, but we do not currently verify or adjust for them. Also, in this work, we used manual prompt creation, where slight variations can significantly affect model outputs Zhao et al. (2021). However, we did not address this variability by testing multiple prompt templates for each prompting technique. Acquiring targeted knowledge is resource-intensive due to the need for manual annotations. Scaling this process is impractical, highlighting the need for automated targeted knowledge acquisition techniques. Not all data points in the dataset include knowledge from ConceptNet, Wikidata, and WordNet due to the incompleteness of these graphs, highlighting the broader challenge of knowledge graph completion.
Acknowledgments
We thank anonymous reviewers for their constructive feedback and Anirudh Govil for his valuable input. This work was supported by NSF grant #2335967: EAGER: Knowledge-guided neurosymbolic AI with guardrails for safe virtual health assistants. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funding organization.
@@ Appendix A Model Details
Model Name WD Knowledge Only CN Knowledge Only WN Knowledge Only All Knowledge Available Falcon 25.14 25.38 25.04 24.14 FlanT5 19.63 28.15 30.70 12.19 GPT2 4.11 2.40 1.00 5.84 Mistral 22.84 22.07 23.34 24.11 Orca 14.62 9.52 9.28 22.24 Zephyr 16.16 12.49 9.75 14.76 CodeT5 21.47 21.64 22.70 20.70 CodeParrot 0.00 0.00 0.00 0.00 GPT-3.5-Turbo 7.13 7.43 7.90 7.20
@@
Figure 6: Prompt Lengths vs. Peak Performance.
Appendix B Implementation Details
We use API requests for GPT-3.5-Turbo and checkpoints from Hugging face [^17] https://huggingface.co/models for open-source models. The models are evaluated with following hyper parameter settings, temperature = 0.1, top_p=0.1 and repetition_penalty=1.2 to elicit more concrete answers for the MCQs. We use Sentence Transformers [^18] https://sbert.net/ to identify semantically similar exemplars and to perform semantic knowledge filtering. We utilize Wikidata knowledge from Wang et al. (2021b), ConceptNet knowledge from conceptnet5191919 https://github.com/commonsense/conceptnet5/wiki/Downloads and WordNet knowledge from Open English WordNet (2023) [^20] https://github.com/globalwordnet/english-wordnet?tab=readme-ov-file.
Appendix C Performance and Additional Results
C.1 Model Performance vs. Prompt Length (PL)
@@ Appendix C Performance and Additional Results
Appendix D Prompts
Figues 7, 8, 9, 10 and 11 illustrates example prompts provided to models.